Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability Problem Solving

Expert replies
by [email protected] » Sat Apr 20, 2013 11:05 pm
Q) A drawer has 2 identical red socks, 4 identical black socks and 6 identical white socks. What is the fewest number of socks that must be pulled out to have at least a 50% chance of pulling out at least one red socks?

A) 1
B) 2
C) 3
D) 4
E) 10
Join the discussion
Source: — Problem Solving |

by vipulgoyal » Sun Apr 21, 2013 12:15 am
my take A
Join the discussion

by neha24 » Sun Apr 21, 2013 12:28 am
i believe that here pair of socks are not to be taken as is usually the case .
so in that case to have fewest number of socks that must be pulled out to have at least a 50% chance of pulling out at least one red socks will be 9 .coz if u take out all the 6 white + 3 black socks then u will be left with 2 identical red and 1 black sock---> in this case the probability of pulling one red sock will be 2/3 which is more than 50 percent but since that option is not there so we can go with 10 hence E
Join the discussion

by Brent@GMATPrepNow » Sun Apr 21, 2013 6:04 am
[email protected] wrote:Q) A drawer has 2 identical red socks, 4 identical black socks and 6 identical white socks. What is the fewest number of socks that must be pulled out to have at least a 50% chance of pulling out at least one red socks?

A) 1
B) 2
C) 3
D) 4
E) 10
So, we have 2 red socks and 10 non-red socks.

To begin, probability questions are great for applying some intuition.

Consider answer choice A. We know that if we draw 1 sock, the probability of selecting a red sock is less than 0.5 (50%) So, eliminate A.

Now consider answer choice E. If we draw 10 socks, the probability of selecting at least one red sock is VERY VERY high. Much greater than 0.5 We can apply some intuition here and conclude that we must be able to reach a probability of 0.5 with few than 10 draws. So, eliminate E.

This leaves us with drawing 2, 3 or 4 socks.

Let's try 3 (answer choice C) and go from there.

If we draw 3 socks, what is P(draw at least 1 red sock)?
When it comes to probability questions involving "at least," it's often a good idea to use the complement.
That is, P(Event A happening) = 1 - P(Event A not happening)
So, here we get: P(getting at least 1 red sock) = 1 - P(not getting at least 1 red sock)
In other words, P(getting at least 1 red sock) = 1 - P(getting zero red socks)

P(getting zero red socks) = P(no red sock on 1st draw AND no red sock on 2nd draw AND no red sock on 3rd draw)
= P(no red sock on 1st draw) x P(no red sock on 2nd draw) x P(no red sock on 3rd draw)
= (10/12) x (9/11) x (8/10)
= 6/11

So, when we select 3 socks, P(at least 1 red sock) = 1 - 6/11
= 5/11 (less than 0.5)
Since we want P(at least 1 red sock) to be 0.5 or greater, we must select more than 3 socks.

IMPORTANT: At this point, the answer must be D, so I'd select D without performing any additional calculations.

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion